# Non-thermal emission from mildly relativistic tidal ejecta of compact objects merger

### v1

@@ Abstract
Mergers of compact objects (binary neutron stars, BNS, or neutron star-black hole, NSBH) with a substantial mass ratio ($q>1.5$) are expected to produce a mildly relativistic ejecta within $\sim 20^{\circ}$ from the orbital plane. We present a semi-analytic approach to calculate the expected synchrotron emission observed from various viewing angles, along with the corresponding radio maps, that are produced by a collisionless shock driven by such ejecta into the interstellar medium. This method is supported by 2D numerical calculations of the full relativistic hydrodynamics with an agreement in the observed emission of $\sim 30\%$. We consider a toroid ejecta with an opening angle of $15^{\circ}\leq\theta_{\text{open}}\leq 30^{\circ}$ and broken power-law mass distribution, $M(>\gamma\beta)\propto(\gamma\beta)^{-s}$ with $s=s_{\rm KN}$ at $\gamma\beta<\gamma_{0}\beta_{0}$ and $s=s_{\rm ft}$ at $\gamma\beta>\gamma_{0}\beta_{0}$ (where $\gamma$ is the Lorentz factor). While the peak flux is dimmer by a factor of $\sim$2-3, and the peak time remains roughly the same (within $20\%$), for various viewing angles compared to isotropic equivalent ejecta ($\theta_{\text{open}}=90^{\circ}$) considered in preceding papers, the radio maps are significantly different from the spherical case. The semi-analytic method can provide information on the ejecta geometry and viewing angle from future radio map observations and, consequently, constrain the ejection mechanism. Due to the larger ejected mass in NSBH merger, this late non-thermal signal can be observed to distances of $\lesssim 200$Mpc for typical parameter values.
@@
pubyear: 2024pagerange: Non-thermal emission from mildly relativistic tidal ejecta of compact objects mergerB
@@ 1 Introduction
Binary mergers involving compact objects with a substantial mass ratio ($q>1.5$) or ones in which a prompt black hole (BH) is produced, such as NSBH or BNS systems, yield expanding ejecta that attains mildly relativistic velocities (for NSBH Rosswog, 2005; Just et al., 2015; Kawaguchi et al., 2015; Kyutoku et al., 2018; Foucart et al., 2017; Foucart, 2020), (for BNS Radice et al., 2018; Bernuzzi et al., 2020; Nedora et al., 2021b), $\gamma\beta>0.1$, through tidal interaction between the BH/NS and the NS (see Bernuzzi, 2020; Kyutoku et al., 2021, for review). Such ejecta is expected to produce non-thermal electromagnetic (EM) emission on a time scale of $\sim$10 years (Nakar & Piran, 2011; Sadeh et al., 2023). BNS and NSBH mergers are likely the sources of (short) gamma-ray bursts (see Mészáros, 2002; Piran, 2004; Nakar, 2007, for reviews). Consequently, these events are expected to produce highly relativistic jets that may dominate the non-thermal emission from the tidally driven ejecta, depending on the jet’s energy, opening angle, and the angle at which it is observed. Typically, because of the jet’s lower mass and higher Lorentz factor, the late-time emission is expected to be dominated by the tidal ejecta. To interpret future observations accurately, it will be necessary to disentangle these two components.
@@
NSBH mergers can exhibit disruptive or non-disruptive behaviors. In the former, the NS is disrupted by the tidal field of the BH, triggering mass ejection, formation of a disk around the BH, and the emission of EM radiation. A disruptive NSBH merger occurs when the disruption radius, $R_{\text{dis}}$, in which the tidal forces of the BH are larger than the gravitational forces that hold the NS as a compact object, is larger than the radius of the innermost stable circular orbit, $R_{\text{ISCO}}$. Conversely, in non-disruptive mergers, the NS plunges as a whole into the BH, and the phenomenon is detectable solely through gravitational wave (GW) measurements. Numerical relativity (NR) simulations of BNS mergers suggest that the mass ratio between the BNS significantly affects the ejecta geometry (Bernuzzi et al., 2020). BNS mergers of comparable mass ($q\sim 1$) eject mass dynamically due to shocks driven by the collision and also due to tidal forces. The ejecta geometry for such mergers is quasi-spherical (Radice et al., 2018; Nedora et al., 2021b). Additional spiral-wave wind component should be considered in case of long-lived neutron star remnant (Nedora et al., 2021b; Radice & Bernuzzi, 2023). In cases where the mass ratio is considerable, $q>1.5$, the ejected mass is concentrated within $20^{\circ}$ from the equatorial plane since most of the mass is ejected by the tidal forces. In BNS mergers in which a prompt BH is produced, the dynamical ejecta is also dominated by the tidal component. Correspondingly, the angular spread of the ejecta beyond the equatorial plane is within $\sim 20^{\circ}$ (Radice et al., 2018; Nedora et al., 2021b). In all cases the NR simulations suggest dynamical ejecta of $\sim$ $10^{-3}-10^{-2}M_{\odot}$ (for various equation of states (EoSs) and mass ratios, see Radice et al., 2020) with velocity profile that is well approximated by the one suggested in Sadeh et al. (2023) and confirmed in Zappa et al. (2023). NSBH mergers with comparable mass have not been observed so far and, thus, have not been considered in NR simulations. NSBH merger simulations of large mass ratio, $q>3$, provide some constraints over the dynamical ejecta of such events (Rosswog, 2005; Just et al., 2015; Kyutoku et al., 2015, 2018; Foucart et al., 2013, 2017; Chen et al., 2024). In cases of tidal disruption, they imply the dynamical ejecta (notice that the spiral wind component (Nedora et al., 2021b; Radice & Bernuzzi, 2023) is absent for NSBH systems) has only a tidal component of $\sim$ $10^{-2}-10^{-1}M_{\odot}$ (for various initial parameters and EoSs). This ejecta shares a similar geometry to BNS dynamical ejecta with a high mass ratio according to numerical relativity simulations (Bernuzzi et al., 2020); furthermore, it is consistent with the velocity profile suggested in Sadeh et al. (2023) for BNS mergers (Kyutoku et al., 2015; Chen et al., 2024). The amount of ejected mass, its geometry, and its velocity distribution depend on the parameters of the merging system (whether it is NSBH or BNS), such as the mass and the spin of the BH (in the case of NSBH), the mass of the NS, the nuclear EoS and the mass ratio, $q$ (Rosswog, 2005; Shibata & Taniguchi, 2008; Lovelace et al., 2013; Kyutoku et al., 2015; Foucart et al., 2017; Hayashi et al., 2022; Chen et al., 2024). For example, for a mass ratio of $q>8$ no considerable mass ejection is expected since the NS is not tidally disrupted (Foucart et al., 2017). Measurements of the EM emission generated by the expanding ejecta will enable us to constrain its properties and, hence, the abovementioned parameters. The opening angle of the ejecta is approximately conserved (Kyutoku et al., 2015) because the velocity direction does not change appreciably once the hydrodynamic merger interaction becomes negligible. The azimuthal opening angle of the ejecta varies from $180^{\circ}$ to $360^{\circ}$ for different initial conditions (Kyutoku et al., 2015; Kawaguchi et al., 2016; Bernuzzi et al., 2020). Margalit & Piran (2015) consider the effect of a non-spherical (and non-relativistic) ejecta structure on the synchrotron light curve. They provide a rough approximation of the peak time scale that is delayed with respect to the isotropic equivalent case and reach the conclusion that the peak flux scale remains roughly the same as in the isotropic equivalent case. However, they do not provide full numerical calculations and base their conclusions on an analytic approximation from Nakar & Piran (2011).
@@
The rate of BNS mergers was estimated by the galactic BNS systems to be $\sim 10\left(\frac{D}{200\text{Mpc}}\right)^{3}\text{yr}^{-1}$ (Phinney, 1991; Pol et al., 2019, 2020). Such estimation is not possible for the rate of NSBH mergers since our galaxy has no known NSBH binary systems. Due to the low number of detected BNS and NSBH mergers via GW, the merger rate obtained from these observations is extremely uncertain. For BNS it is between $0.1$ to $14\left(\frac{D}{200\text{Mpc}}\right)^{3}\text{yr}^{-1}$ and for NSBH merger it is between $0.1$ to $1.5\left(\frac{D}{200\text{Mpc}}\right)^{3}\text{yr}^{-1}$ (LIGO Scientific Collaboration et al., 2023; The LIGO Scientific Collaboration et al., 2024). The fraction of EM bright events out of the detectable (via GW) NSBH mergers events is currently unknown due to the low statistics (only three confirmed events so far in the third and fourth LIGO-Virgo-KAGRA run (Abbott et al., 2021; The LIGO Scientific Collaboration et al., 2024). However, first attempts have been made (Fragione, 2021; Biscoveanu et al., 2023; The LIGO Scientific Collaboration et al., 2024), suggesting a fraction of $\sim 10-30\%$. Martineau et al. (2024) estimated the ejected mass in the observed event, GW230529, to be negligible, $\sim 10^{-3}M_{\odot}-10^{-5}M_{\odot}$, due to the low BH spin $\chi_{\text{BH}}\lesssim 0.1$. While the mass ratio in NSBH mergers is likely to be above $q=1.5$ (considered the observed mergers so far), for the observed BNS mergers GW170817 and GW190425, the estimation varies between $q\approx 1$ to $q\approx 2.5$ for different spin priors.
@@
This paper explores synchrotron emission spanning the radio-to-X-ray range, originating from a collisionless shock driven by mildly relativistic toroidal ejecta of an opening angle of between $15^{\circ}$ to $30^{\circ}$ and a broken power-law dependence of mass on momentum, with parameter values characteristic of the results of numerical calculations of the BNS ejecta (see §2). The ejecta expands into a uniform interstellar medium (ISM) characterized by a (number) density $n$. We derive a semi-analytic calculation method for the case of axisymmetric distribution of mass (as in Gompertz et al., 2023) exhibiting a power-law momentum-mass dependence (Sadeh et al., 2023), where the parameter values are consistent with the results of numerical calculations of the tidal ejecta (Kyutoku et al., 2015, 2018; Foucart et al., 2017; Bernuzzi et al., 2020; Nedora et al., 2021b). This semi-analytic method is valid for as long as the reverse shock did not cross the fast tail part of the ejecta, during this phase the shocked plasma is not expected to go through a significant lateral expansion, so we assume angular independent distribution of shocked plasma parameters. We derive the non-thermal emission under the assumption of a power-law energy distribution for electrons, given by $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. Here, $\gamma_{e}$ represents the electron Lorentz factor within the plasma’s rest frame, while $p$ is within the range of $2\leq p\leq 2.5$. Concurrently, fractions $\varepsilon_{e}$ and $\varepsilon_{B}$ correspond to the post-shock internal energy density proportions attributed to non-thermal electrons and magnetic fields, respectively. This phenomenological description, capturing the post-shock plasma conditions, finds support across a diverse spectrum of observations and plasma calculations encompassing both relativistic and non-relativistic shocks (see Blandford & Eichler, 1987; Keshet & Waxman, 2005; Waxman, 2006; Keshet, 2006; Bykov & Treumann, 2011; Sironi et al., 2013; Pohl et al., 2020; Ligorini et al., 2021; Kobzar et al., 2021). We verified the accuracy of the semi-analytic results for the non-thermal emission presented in this paper through a comprehensive comparison with full 2D numerical computations across a wide range of parameter configurations.
@@
The paper is organized as follows. In § 2, we describe the framework of the calculations presented later on. We derive the semi-analytic calculation for the synchrotron light curve and corresponding radio map in § 3. Then, we describe the 2D relativistic-hydrodynamics numerical calculation and the post-processing we use to produce light curves in § 4. The accuracy of the semi-analytic method is estimated by comparison to the detailed numeric calculations in § 5, while in § 6, we provide a few examples for light curves and radio maps produced by the semi-analytic approach. Finally, our conclusions are summarized in § 7.
@@ 2.1 Geometry description
In Sadeh et al. (2023, 2024), we considered a spherical ejecta with the following mass profile
@@
with parameter values characteristic of the results of numerical calculations of the BNS ejecta; $0.50.3<\beta_{0}<$ 0.5, $125<s_{\text{ft}}<$ 12, $30.5<s_{\text{KN}}<$ 3, and $10^{-6}<M_{0}<10^{-4}$. This analytic form provides a good approximation for the variety of ejecta profiles obtained in NR simulations of BNS mergers (Zappa et al., 2023). A merger of compact objects with a mass ratio of $q>1.5$, or one in which a prompt BH is produced, is expected to eject mass mainly within an angle of $\sim 20^{\circ}$ from the equatorial plane and follow the same velocity profile as the dynamical ejecta of merger with comparable masses (Kyutoku et al., 2015; Foucart et al., 2017; Radice et al., 2018; Bernuzzi et al., 2020; Nedora et al., 2021b). The ejecta azimuthal structure varies from full azimuthal symmetry to a crescent shape with an azimuthal opening of $\phi\approx 180^{\circ}$ (Kyutoku et al., 2015). The geometric structure we consider to describe such ejecta is a toroid (sphere truncated by a double-sided cone), which limits us to the case of full azimuthal symmetry, with a mass profile that is given by Eq. 1, and various opening angles that are consistent with the results of NR simulations ($15^{\circ}-30^{\circ}$), at some initial lab time $t_{0}$ (rest frame of the ISM). The angle between the equatorial plane and the ejecta "edge" is defined as $\theta_{\text{open}}$, while $\theta_{\text{v}}$ is the angle between the symmetry axis and the line of sight. In Fig. 1, we provide a sketch illustrating the ejecta geometry. As the ejecta expands, a forward shock is driven into the ISM, and a reverse shock is driven into the ejecta.

### v2

# Late-time non-thermal emission from mildly relativistic tidal ejecta of compact objects merger

@@ Abstract
Mergers of compact objects (binary neutron stars, BNS, or neutron star-black hole, NSBH) with a substantial mass ratio ($q>1.5$) are expected to produce a mildly relativistic ejecta within $\sim 20^{\circ}$ from the equatorial plane. We present a semi-analytic approach to calculate the expected synchrotron emission observed from various viewing angles, along with the corresponding radio maps, that are produced by a collisionless shock driven by such ejecta into the interstellar medium. This method reproduces well (up to $\sim 30\%$ deviations) the observed emission produced by 2D numerical calculations of the full relativistic hydrodynamics. We consider a toroidal ejecta with an opening angle of $15^{\circ}\leq\theta_{\text{open}}\leq 30^{\circ}$ and broken power-law mass distribution, $M(>\gamma\beta)\propto(\gamma\beta)^{-s}$ with $s=s_{\rm KN}$ at $\gamma\beta<\gamma_{0}\beta_{0}$ and $s=s_{\rm ft}$ at $\gamma\beta>\gamma_{0}\beta_{0}$ (where $\gamma$ is the Lorentz factor). The parameter values are chosen to characterize merger calculation results- a "shallow" mass distribution, $31<s_{\rm KN}<$ 3, for the bulk of the ejecta (at $\gamma\beta\approx 0.2$), and a steep, $s_{\rm ft}>5$, "fast tail" mass distribution. While the peak flux is dimmer by a factor of $\sim$2-3, and the peak time remains roughly the same (within $20\%$), for various viewing angles compared to isotropic equivalent ejecta ($\theta_{\text{open}}=90^{\circ}$) considered in preceding papers, the radio maps are significantly different from the spherical case. The semi-analytic method can provide information on the ejecta geometry and viewing angle from future radio map observations and, consequently, constrain the ejection mechanism. For NSBH mergers with a significant mass ejection ($\sim 0.1M_{\odot}$), this late non-thermal signal can be observed to distances of $\lesssim 200$Mpc for typical parameter values.
@@
pubyear: 2024pagerange: Late-time non-thermal emission from mildly relativistic tidal ejecta of compact objects mergerB
@@ 1 Introduction
Binary mergers involving compact objects with a substantial mass ratio ($q>1.5$) or ones in which a prompt black hole (BH) is produced, such as NSBH (Rosswog, 2005; Just et al., 2015; Kawaguchi et al., 2015; Kyutoku et al., 2018; Foucart et al., 2017; Foucart, 2020) or BNS (Radice et al., 2018; Bernuzzi et al., 2020; Nedora et al., 2021b) systems, yield expanding ejecta that attains mildly relativistic velocities, $\gamma\beta>0.1$, through tidal interaction between the BH/NS and the NS (see Bernuzzi, 2020; Kyutoku et al., 2021; Chen et al., 2024, for review). Such ejecta is expected to produce non-thermal electromagnetic (EM) emission on a time scale of $\sim$10 years (Nakar & Piran, 2011; Sadeh et al., 2023). BNS and NSBH mergers are likely the sources of (short) gamma-ray bursts (see Mészáros, 2002; Piran, 2004; Nakar, 2007, for reviews). Consequently, these events are expected to produce highly relativistic jets that may dominate the non-thermal emission from the tidally driven ejecta, depending on the jet’s energy, opening angle, and the angle at which it is observed. Typically, because of the jet’s lower mass and higher Lorentz factor, the late-time emission is expected to be dominated by the tidal ejecta. To interpret future observations accurately, it will be necessary to disentangle these two components.
@@
NSBH mergers can exhibit disruptive or non-disruptive behaviors. In the former, the NS is disrupted by the tidal field of the BH, triggering mass ejection, formation of a disk around the BH, and the emission of EM radiation. A disruptive NSBH merger occurs when the disruption radius, $R_{\text{dis}}$, at which the tidal forces of the BH are larger than the gravitational forces that hold the NS as a compact object, is larger than the radius of the innermost stable circular orbit, $R_{\text{ISCO}}$. Conversely, in non-disruptive mergers, the NS plunges as a whole into the BH, and the phenomenon is detectable solely through gravitational wave (GW) measurements. Numerical relativity (NR) simulations of BNS mergers suggest that the mass ratio between the BNS significantly affects the ejecta geometry (Bernuzzi et al., 2020; Bernuzzi et al., 2024). BNS mergers of comparable mass ($q\sim 1$) eject mass dynamically due to shocks driven by the collision and also due to tidal forces. The ejecta geometry for such mergers is quasi-spherical (Radice et al., 2018; Nedora et al., 2021b). Additional spiral-wave wind component (Nedora et al., 2021b; Radice & Bernuzzi, 2023) should be considered in case of long-lived neutron star remnant (absent for NSBH systems). In cases where the mass ratio is considerable, $q>1.5$, the ejected mass is concentrated within $20^{\circ}$ from the equatorial plane since most of the mass is ejected by the tidal forces. In BNS mergers in which a prompt BH is produced, the dynamical ejecta is also dominated by the tidal component. Correspondingly, the angular spread of the ejecta beyond the equatorial plane is within $\sim 20^{\circ}$ (Radice et al., 2018; Nedora et al., 2021b). In all cases the NR simulations suggest dynamical ejecta of $\sim$ $10^{-3}-10^{-2}M_{\odot}$, for various equation of states (EoSs) and mass ratios (see Radice et al., 2020), with velocity profile that is well approximated by the one suggested in Sadeh et al. (2023) and confirmed in Zappa et al. (2023). NSBH mergers with comparable mass have not been observed so far and, thus, have not been considered in NR simulations. NSBH merger simulations of large mass ratio, $q>3$, provide some constraints over the dynamical ejecta of such events (Rosswog, 2005; Just et al., 2015; Kyutoku et al., 2015, 2018; Foucart et al., 2013, 2017; Chen et al., 2024). In cases of tidal disruption, they imply the dynamical ejecta has only a tidal component of $\sim$ $10^{-2}-10^{-1}M_{\odot}$, for various initial parameters and EoSs. This ejecta shares a similar geometry to BNS dynamical ejecta with a high mass ratio according to numerical relativity simulations (Bernuzzi et al., 2020; Bernuzzi et al., 2024); furthermore, it is consistent with the velocity profile suggested in Sadeh et al. (2023) for BNS mergers (Kyutoku et al., 2015; Chen et al., 2024). The amount of ejected mass, its geometry, and its velocity distribution depend on the parameters of the merging system (whether it is NSBH or BNS), such as the mass and the spin of the BH (in the case of NSBH), the mass of the NS, the nuclear EoS and the mass ratio, $q$ (Rosswog, 2005; Shibata & Taniguchi, 2008; Lovelace et al., 2013; Kyutoku et al., 2015; Foucart et al., 2017; Hayashi et al., 2022; Chen et al., 2024). For example, for a mass ratio of $q>8$, no considerable mass ejection is expected since the NS is not tidally disrupted (Foucart et al., 2017). Measurements of the EM emission generated by the expanding ejecta will enable us to constrain its properties and, hence, the abovementioned parameters. The opening angle of the ejecta is approximately conserved (Kyutoku et al., 2015) because the velocity direction does not change appreciably once the hydrodynamic merger interaction becomes negligible. The azimuthal opening angle of the ejecta varies from $180^{\circ}$ to $360^{\circ}$ for different initial conditions (Kyutoku et al., 2015; Kawaguchi et al., 2016; Bernuzzi et al., 2020). Margalit & Piran (2015) consider the effect of a non-spherical (and non-relativistic) ejecta structure on the synchrotron light curve. They provide a rough approximation of the peak time scale that is delayed with respect to the isotropic equivalent case and reach the conclusion that the peak flux scale remains roughly the same as in the isotropic equivalent case. However, they do not provide full numerical calculations and base their conclusions on an analytic approximation from Nakar & Piran (2011). It should be noted here that more material is expected to be ejected from the remnant disk (Fernández et al., 2017). Due to the lower velocities and time delay compared to the dynamical ejecta, no mixing between the two components is expected (Fernández et al., 2015). Furthermore, since the non-thermal signal is highly sensitive to the shocked plasma Lorentz factor (Sadeh et al., 2023), the lower velocity of the disk outflow (Fernández et al., 2017, 2019) would make its potential non-thermal signal much less bright.
@@
The rate of BNS mergers was estimated by the Galactic BNS systems to be $\sim 10\left(\frac{D}{200\text{Mpc}}\right)^{3}\text{yr}^{-1}$ (Phinney, 1991; Pol et al., 2019, 2020). Such estimation is not possible for the rate of NSBH mergers since our Galaxy has no known NSBH binary systems. Due to the low number of detected BNS and NSBH mergers via GW, the merger rate obtained from these observations is extremely uncertain. For BNS mergers it is between $0.1$ to $14\left(\frac{D}{200\text{Mpc}}\right)^{3}\text{yr}^{-1}$ and for NSBH mergers it is between $0.1$ to $1.5\left(\frac{D}{200\text{Mpc}}\right)^{3}\text{yr}^{-1}$ (LIGO Scientific Collaboration et al., 2023; The LIGO Scientific Collaboration et al., 2024). The fraction of EM bright events out of the GW detectable NSBH mergers events is currently unknown due to the low statistics- only three confirmed events so far in the third and fourth LIGO-Virgo-KAGRA run (Abbott et al., 2021; The LIGO Scientific Collaboration et al., 2024). However, first attempts have been made (Fragione, 2021; Biscoveanu et al., 2023; The LIGO Scientific Collaboration et al., 2024), suggesting a fraction of $\sim 10-30\%$. Martineau et al. (2024) estimated the ejected mass in the observed event, GW230529, to be negligible, $\sim 10^{-5}M_{\odot}-10^{-3}M_{\odot}$, due to the low BH spin $\chi_{\text{BH}}\lesssim 0.1$. While the mass ratio in NSBH mergers is likely to be above $q=1.5$ considered the observed mergers so far, for the observed BNS mergers GW170817 and GW190425, the mass ratio estimate varies between $q\approx 1$ to $q\approx 2.5$ for different spin priors.
@@
This paper explores synchrotron emission spanning the radio-to-X-ray range, originating from a collisionless shock driven by mildly relativistic toroidal ejecta with an opening angle of $15^{\circ}$ to $30^{\circ}$ and a broken power-law dependence of mass on momentum, with parameter values characteristic of the results of numerical calculations of the BNS ejecta (see §2). The ejecta expands into a uniform interstellar medium (ISM) characterized by a number density $n$. We derive a semi-analytic calculation method for the case of axisymmetric distribution of mass (as in Gompertz et al., 2023) exhibiting a power-law momentum-mass dependence (Sadeh et al., 2023), where the parameter values are consistent with the results of numerical calculations of the tidal ejecta (Kyutoku et al., 2015, 2018; Foucart et al., 2017; Bernuzzi et al., 2020; Nedora et al., 2021b; Chen et al., 2024; Bernuzzi et al., 2024). This semi-analytic method is valid for as long as the reverse shock did not cross the fast tail part of the ejecta, during this phase the shocked plasma is not expected to go through a significant lateral expansion, so we assume angular independent distribution of shocked plasma parameters. We derive the non-thermal emission spectra under the assumption of a power-law energy distribution for electrons, given by $dn_{e}/d\gamma_{e}\propto\gamma_{e}^{-p}$. Here, $\gamma_{e}$ represents the electron Lorentz factor within the plasma’s rest frame, where $p$ is within the range of $2\leq p\leq 2.5$. Concurrently, the fractions $\varepsilon_{e}$ and $\varepsilon_{B}$ correspond to the post-shock internal energy density proportions attributed to non-thermal electrons and magnetic fields, respectively. This phenomenological description, capturing the post-shock plasma conditions, finds support across a diverse spectrum of observations and plasma calculations encompassing both relativistic (see Keshet & Waxman, 2005; Waxman, 2006; Keshet, 2006; Bykov & Treumann, 2011; Sironi et al., 2013; Kobzar et al., 2021) and non-relativistic shocks (see Blandford & Eichler, 1987; Pohl et al., 2020; Ligorini et al., 2021). We verified the accuracy of the semi-analytic results for the non-thermal emission presented in this paper through a comprehensive comparison with full 2D numerical computations across a wide range of parameter configurations.
@@
The paper is organized as follows. In § 2, we describe the framework of the calculations presented later on. We derive the semi-analytic calculation for the synchrotron light curve and corresponding radio map in § 3. Then, we describe the 2D relativistic-hydrodynamics numerical calculation and the post-processing we use to produce light curves in § 4. The accuracy of the semi-analytic method is estimated by comparison with the detailed numeric calculations in § 5, while in § 6, we provide a few examples for light curves and radio maps produced by the semi-analytic approach. Finally, our conclusions are summarized in § 7.
@@ 2.1 Ejecta geometry description
In Sadeh et al. (2023); Sadeh & Waxman (2024), we considered a spherical ejecta with the following mass profile
@@
with parameter values characteristic of the results of numerical calculations of the BNS ejecta; $0.50.3<\beta_{0}<$ 0.5, $125<s_{\text{ft}}<$ 12, $30.5<s_{\text{KN}}<$ 3, and $10^{-6}M_{\odot}<M_{0}<10^{-4}M_{\odot}$. This analytic form provides a good approximation for the variety of ejecta profiles obtained in NR simulations of BNS mergers (Zappa et al., 2023). A merger of compact objects with a mass ratio of $q>1.5$, or one in which a prompt BH is produced, is expected to eject mass mainly within an angle of $\sim 20^{\circ}$ from the equatorial plane and follow the same velocity profile as the dynamical ejecta of mergers with comparable masses (Kyutoku et al., 2015; Foucart et al., 2017; Radice et al., 2018; Bernuzzi et al., 2020; Nedora et al., 2021b; Bernuzzi et al., 2024). The ejecta azimuthal structure varies from full azimuthal symmetry to a crescent shape with an azimuthal opening of $\phi\approx 180^{\circ}$ (Kyutoku et al., 2015). The geometric structure we consider to describe such ejecta is a toroid (sphere truncated by a double-sided cone), which limits us to the case of full azimuthal symmetry, with a mass profile that is given by Eq. 1, and various opening angles, $15^{\circ}-30^{\circ}$, that are consistent with the results of NR simulations, at some initial lab time $t_{0}$ (rest frame of the ISM). The angle between the equatorial plane and the ejecta "edge" is defined as $\theta_{\text{open}}$, while $\theta_{\text{v}}$ is the angle between the symmetry axis and the line of sight. In Fig. 1, we provide a sketch illustrating the ejecta geometry. As the ejecta expands, a forward shock is driven into the ISM, and a reverse shock is driven into the ejecta.